1. Sphere 1 has surface area A1 and volume V1 and sphere 2 has surface area A2 and volume V2. If the radius of sphere 2 is 3.86 times the radius of sphere 1, what is the ratio of the areas A2/A1?


2. What is the ratio of the volumes V2/V1
Answer #1

We know that area of a sphere is given as

A = 4r2

given

r2 = 3.86 * r1

so,

A2 / A1 = r22 / r12

4 will get cancelled

A2 / A1 = 3.862 * r12 / r12

A2 / A1 = 14.89

----------------------------------------------------------------

volume of sphere is given as

V = 4/3r3

4/3 will get cancelled, we get

V2 / V1 = r23 / r13

V2/ V1 = 3.863

V2 / V1 = 57.51